If the population size is equal to 3 and the sample size is equal to 5. How many possible samples can be drawn using the formula of with replacement?




1
Expert's answer
2022-05-24T23:04:26-0400

The number of possible samples which can be selected with replacement is

"N^n=3^5=243."